# Data-Driven Computation of Polytopic Invariant Sets for Noisy Nonlinear Systems

### v1

@@
Abstract.
This paper presents a data-driven framework for computing robust, convex polytopic contractive set for constrained noisy nonlinear systems where an analytical model is not available. Our approach utilizes a finite set of collected noisy measurements to construct a polytopic set that bounds all possible system parameters compatible to available information. Based on previous results, we contribute to provide a sufficient condition for a set to be contractive using Difference of Convex functions for a noisy nonlinear system, while the model is not available. Robustness with respect to unknown model of system is guaranteed by requiring that the computed contractive set is invariant for all possible system models that are compatible with the noisy measurements. We present a tractable, optimization-based algorithm that implements this condition to compute the largest possible contractive set within the state constraint set for the unknown, noisy nonlinear system which is subjected to both state and input constraints. The effectiveness of the proposed methodology is demonstrated with a numerical example.
footnotetext: Sahand Kiani is a PhD student in the Department of Electrical and Computer Engineering, Pennsylvania State University, State College, PA 16801, USA. Constantino M. Lagoa is faculty with the Department of Electrical and Computer Engineering, Pennsylvania State University, State College, PA 16801, USA.
1. Introduction
@@
One of the interesting aspects in the study of dynamical systems is the analysis of their invariant sets. The task of computing such a set is of significant practical importance, as a contractive set not only guarantees invariance but also associated with the requirements of safety and stability of systems [blanchini1999set]. This property makes these sets essential for designing high-performance controllers, particularly for their use as terminal sets in Model Predictive Control to ensure recursive feasibility and stability. While the computation of invariant sets is a well-established field for linear systems, significant challenges persist for nonlinear systems, especially when an accurate mathematical model is unavailable and the dynamics are affected by uncertainty. Developing computationally tractable methods to find such sets under these conditions is a key objective in modern control research.
The computation of invariant sets can impose significant challenges when the system of interest is not linear, particularly in the presence of uncertainty. Even when a precise mathematical model is known, the inherent nonlinearity often results in sets with complex, non-convex geometries that are difficult to represent, and the requirement for robustness against all possible disturbances adds another layer of complexity. These challenges are significantly magnified in the data-driven setting, where an explicit model of the system is unavailable, rendering the vast majority of traditional analysis tools inapplicable.
1.1. Related Work
@@
The computation of invariant sets for nonlinear systems has been addressed using various optimization-based techniques. For instance, the authors in [10123035] propose a method for perturbed nonlinear sampled-data systems by formulating the problem as a non-convex program. They introduce a tailored successive convexification algorithm that leverages a zonotopic representation for the sets and is designed to find a large region of safe operation by ensuring the sequence of candidate sets is monotonically increasing in volume. Other approaches for handling constrained nonlinear systems focus on approximating the one-step reachable set. For example, the work in [bravo2005computation] presents a method to compute a guaranteed inner approximation of this set for known nonlinear models. The authors develop a specialized branch and bound algorithm based on interval arithmetic, which is followed by a second procedure to find a simple inner polytopic representation of the computed region.
@@
An alternative line of research approaches the computation of invariant sets using graph-based methods. The authors in [decardi2021computing], for example, cast the problem of finding the largest robust control invariant sets for constrained nonlinear systems as a problem in graph theory. Their methodology involves approximating the system dynamics with a directed graph and then applying an iterative subdivision algorithm to generate a sequence of outer approximations that is proven to converge to the true largest robust control invariant sets.
@@
The literature on computing robust invariant sets for nonlinear systems features several prominent model-based strategies, including optimization-based methods using successive convexification, set-approximation techniques using interval arithmetic, and graph-based approaches that rely on state-space discretization. A unifying limitation of these works is their dependence on an explicit mathematical model, which motivates the purely data-driven framework proposed in this paper.
1.2. Key Contribution
@@
The primary contribution of this paper is the development of a data-driven framework for computing robust convex polytopic contractive sets for a class of unknown, noisy nonlinear systems. We consider a unknown, noisy nonlinear system which is subject to both state and input constraints. Through the paper, it is assumed that the upper bound of the process noise is known. Our approach begins by constructing a polytopic consistency set determined from a set of collected noisy measurements. The core theoretical contribution of this work is the adaptation of an established sufficient condition for contractiveness, originally developed for deterministic noisy nonlinear systems, to our data-driven framework. This will be achieved by ensuring that the computed set is contractive for all systems compatible with available information. Finally, we present a tractable, optimization-based algorithm that operationalizes our sufficient condition to compute the largest possible $\lambda$-contractive set within the state constraint set by solving a convex optimization problem.
@@
Notation. The set of real numbers is denoted by $\mathbb{R}$, and the $n$-dimensional Euclidean space is $\mathbb{R}^{n}$.The state vector at time step $k$ is denoted by $x_{k}\in\mathbb{R}^{n}$, with its $j$-th component denoted by $x_{j,k}$. The operator $\nabla_{x}$ denotes the gradient with respect to the vector $x$. The infinity norm of a vector is denoted by $\|\cdot\|_{\infty}$. For every $n\in\mathbb{N}$, define $N_{n}=\{x\in\mathbb{N}:1\leq x\leq n\}$. The support function of a set $\Gamma$ is $\phi_{\Gamma}(\cdot)$. The vector of all ones with the proper dimension is denoted by $\mathbf{1}$. The interior of a set is denoted by $\text{int}(\cdot)$. The notation $\lambda\Omega$ represents the scaling of the set $\Omega$ by the constant scalar $\lambda$. Consider The matrix $H\in\mathbb{R}^{n\times n}$ The notation $H_{i}$ represents the $i$-th row of this matrix. Throughout this paper, the notation $\theta=\text{vec}(\Theta^{T})$ defines the vector $\theta$ as the vectorization of the transpose of the matrix $\Theta$. Given a set $P$, co($P$) denotes its convex hull,
Paper Organization: The remainder of this paper is organized as follows. Section 2 reviews the necessary preliminaries on set invariance and the properties of DC functions. In Section 3, we provide a formal statement of the data-driven problem. Our main theoretical results, including the data-driven sufficient condition and the proposed computational algorithm, are presented in Section 4. Section 5 validates the effectiveness of our approach with a numerical example. Finally, Section 6 concludes the paper and suggests directions for future work.
2. Preliminary Results
@@
Consider the discrete-time nonlinear system as follows
(1) $x_{k+1}=f(x_{k},u_{k})+w_{k},$
@@
The sets $X$, $U$, and $W$ are assumed to be compact, convex, and to contain the origin in their respective interiors.
2.1. Difference of Convex Functions
A key advantage of the DC framework is its generality, as a wide class of nonlinear functions can be represented as the difference of two convex functions. This structure is particularly useful because it allows for the application of powerful tools from convex analysis to otherwise non-convex problems. We therefore make the following assumption on the nonlinear dynamics.
@@ 2.1. Difference of Convex Functions
The system dynamics, represented by $f(x,u):X\times U\to\mathbb{R}^{n}$, is assumed to be a vector of DC functions, differentiable at $(0,0)$. This implies there exist two functions $g_{j},h_{j}:X\times U\to\mathbb{R}$ for all $j\in N_{n}$ of vector $f(x,u)$, which are convex $\forall(x_{k},u_{k})\in X\times U$, that satisfy the following conditions:
@@
The class of DC functions is quite large: Every function with continuous second partial derivatives is a DC function on a convex compact set of its domain [horst1999dc]. Consequently, the requirement in Assumption 2 that the system dynamics be a DC function is not a significant restriction, ensuring that the methods proposed in this paper are applicable to a broad class of nonlinear systems.
With this functional framework established, we now introduce the key definitions of $\lambda$-contractiveness and sufficient condition for a set to be contractive for nonlinear systems that are fundamental to our analysis.
$2.$ $\lambda$-Contractiveness for Nonlinear Systems
Before presenting our main data-driven results, it is necessary to review the key theoretical concepts from set-theoretic control upon which our work is built. Therefore, we start by the following definition:
@@
[272351]: A convex compact set $\Omega$ with $0\in\text{int}(\Omega)$ is $\lambda$-contractive for system $x_{k+1}=f(x_{k},u_{k})+w_{k}$ and constraints $x_{k}\in X$ and $u_{k}\in U$ if $\Omega\subseteq X$ and for every $x_{k}\in\Omega$ there exists a $u_{k}\in U$ such that $f(x_{k},u_{k})+w_{k}\in\lambda\Omega$, $\forall w_{k}\in W$, with contraction factor $\lambda\in[0,1]$.
@@
In the specific case where the set $\Omega$ is a polytope described by a system of linear inequalities, i.e., $\Omega=\{x\in\mathbb{R}^{n}:Hx\leq\textbf{1}\}$, the condition for $\lambda$-contractiveness can be stated more directly. For every $x_{k}\in\Omega$, there exist a control input $u_{k}\in U$ such that $\forall w_{k}\in W$, the resulting $f(x_{k},u_{k})+w_{k}$ satisfies the polytope’s defining inequalities scaled by the factor $\lambda$:
$H(f(x_{k},u_{k})+w_{k})\leq\lambda\mathbf{1},\quad\forall x_{k}\in\Omega,\forall w_{k}\in W.$
Building on the definition of a robust control invariant set, we now review a key result from [fiacchini2010computation]. The main contribution of that work was to leverage the properties of DC functions to establish a sufficient condition for a set to be $\lambda$-contractive for nonlinear systems.
@@
[boyd2004convex]: Given a set $\Gamma\subseteq\mathbb{R}^{n}$, the support function of $\Gamma$ evaluated at $c\in\mathbb{R}^{n}$ is defined as: $\phi_{\Gamma}(c)=\sup_{x\in\Gamma}c^{T}x$.
@@
[fiacchini2010computation]: Let Assumptions 1 and 2 hold. Given the vector of DC function $f(\cdot,\cdot):X\times U\to\mathbb{R}^{n}$, which its components are decomposed as $f_{j}(x,u)=g_{j}(x,u)-h_{j}(x,u),\forall j\in N_{n}$ with convex functions $g_{j}(x,u)$ and $h_{j}(x,u)$, and $c\in\mathbb{R}^{n}$, define $F(x,u,c):X\times U\times\mathbb{R}^{n}\to\mathbb{R}$ as the function
(2) $F(x,u,c)=$ $\sum_{j\in b_{+}}c_{j}(g_{j}(x,u)-h_{j}^{L}(x,u))$
@@
It can be shown that for any $c\in\mathbb{R}^{n}$ inequality $c^{T}f(x,u)\leq F(x,u,c)$ holds $\forall(x,u)\in X\times U$[fiacchini2010computation]. Another important property of function $F(x,u,c)$ is reviewed in the following property, which the reader is referred to the paper for proofs [fiacchini2010computation].
@@
[fiacchini2010computation]: Let Assumptions 1 and 2 hold. Given the DC function $f(\cdot,\cdot):X\times U\to\mathbb{R}^{n}$ as in (1), then for every $c\in\mathbb{R}^{n}$, the function $F(\cdot,\cdot,c)$ defined in (2) is convex in $(x,u)\in X\times U$.
@@
[fiacchini2010computation]: Let Assumptions 1 and 2 hold. Consider a polytope $\Omega=\{x\in\mathbb{R}^{n}:Hx\leq\mathbf{1}\}\subseteq X$, and DC system (1) which is subject to process noise. If there exist control actions defined at the vertices $u^{p}=u_{v}^{p}\in U$, $\forall p\in\mathbb{N}_{n_{v}}$, such that
(3) $F(v^{p},u_{v}^{p},H_{i}^{T})\leq\lambda_{w}-\phi_{W}(H_{i}^{T}),\quad\forall p\in\mathbb{N}_{n_{v}},\forall i\in\mathbb{N}_{n_{h}},$
where $\lambda_{w}\in[0,1]$ represents the contraction factor for noisy nonlinear system, then $\Omega$ is a $\lambda$-contractive set for system (1) and constraints $x\in X,u\in U$.
The preceding review establishes the theoretical groundwork for model-based control invariance. Building upon these concepts, we are now going to formally state the data-driven problem that is the focus of this work.
3. Problem Statement
This paper develops a data-driven method for computing $\lambda$-contractive sets for the noisy nonlinear system (1), utilizing only a finite set of measured state-input data. The proposed approach relies on the following assumptions:
@@ 3. Problem Statement
The unknown nonlinear function $f(x_{k},u_{k})$ is assumed to lie in the span of a known dictionary of $d_{f}$ basis functions, $\phi(x,u):X\times U\to\mathbb{R}^{d_{f}}$. Consequently, there exists an unknown parameter matrix $\Theta\in\mathbb{R}^{n\times d_{f}}$ such that the system dynamics can be expressed as $f(x_{k},u_{k})=\Theta\phi(x_{k},u_{k})$.
In this context, the problem under consideration in this paper can be formally stated as:
@@
Under Assumptions 13, the objective is to compute the largest polytopic $\lambda$-contractive set, $\Omega\subseteq X$, using the finite dataset $\mathcal{D}=\left\{x_{k},u_{k},x_{k+1}\}_{k=0}^{T-1}\right$.. The computed set must be $\lambda$-contractive for all possible system dynamics that are consistent with the data dictionary, $\mathcal{D}$, and the known noise upper bound $\epsilon$.
4. Data-Driven Computation of Robust Polytopic Control Invariant Set
This section presents our main contribution: a data-driven algorithm for the computation of a robust polytopic $\lambda$-contractive set.

### v2

# Robust Data-Driven Invariant Sets for Nonlinear Systems

@@
Abstract
The synthesis of robust invariant sets for nonlinear systems has traditionally been hindered by the inherent non-convexity and a strict reliance on exact analytical models. This paper presents a purely data-driven framework to compute robust polytopic contractive sets for unknown nonlinear systems operating under persistent bounded process noise and state-input constraints. Rather than attempting to identify a single, potentially nominal model, we utilize a finite data set to construct a polytopic consistency set—a rigorous geometric boundary encapsulating all possible system dynamics compatible with the noisy measurements. The core contribution of this work extends an established sufficient condition for $\lambda$-contractiveness into the data-driven setting. Crucially, we prove that enforcing this condition strictly over the vertices of the consistency set guarantees robust invariance.
I INTRODUCTION
@@
One of the aspects in the study of dynamical systems is the analysis of their invariant sets [3]. The task of computing such a set is of significant practical importance, as these sets not only guarantee invariance but are also associated with the requirements of safety and stability of systems in many applications [11, 10]. This property makes these sets essential for designing high-performance controllers, particularly for their use as terminal sets in Model Predictive Control to ensure recursive feasibility and stability. Approximating invariant sets for nonlinear systems is significantly more complex than for their linear counterparts, primarily due to the non-convex nature of the resulting sets and the requirement for robustness against uncertainty. These challenges are magnified in the data-driven setting, where the lack of an explicit model invalidates most traditional analysis tools.
I-A Related Work
@@
The computation of invariant sets for nonlinear systems has been addressed using various optimization-based techniques. For instance, the authors in [13] propose a method for perturbed nonlinear sampled-data systems by formulating the problem as a non-convex program. They introduce a tailored successive convexification algorithm that leverages a zonotopic representation for the sets and is designed to find a large region of safe operation by ensuring the sequence of candidate sets is monotonically increasing in volume. Other approaches for handling constrained nonlinear systems focus on approximating the one-step reachable set. For example, the work in [5] presents a method to compute a guaranteed inner approximation of the set for known nonlinear models. The authors develop a specialized branch and bound algorithm based on interval arithmetic, which is followed by a second procedure to find a simple inner polytopic representation of the computed region.
@@
An alternative line of research approaches the computation of invariant sets using graph-based methods. The authors in [7], for example, cast the problem of finding the largest robust control invariant sets for constrained nonlinear systems as a problem in graph theory. Their methodology involves approximating the system dynamics with a directed graph and then applying an iterative subdivision algorithm to generate a sequence of outer approximations that is proven to converge to the true largest robust control invariant set.
@@
Existing methods for computing robust invariant sets, such as successive convexification, interval arithmetic, and graph-based discretization, are fundamentally limited by their reliance on an explicit mathematical model. This limitation motivates the purely data-driven framework proposed in this paper.
I-B Key Contribution
@@
The primary contribution of this paper is the development of a data-driven framework for approximating a robust convex polytopic contractive set for a class of noisy nonlinear systems. We consider a class of uncertain nonlinear systems characterized by known polynomial basis functions with unknown parameters. The system is subject to persistent, bounded process noise as well as both state and input constraints. The core theoretical contribution is the extension of an established sufficient condition for contractiveness—originally developed for known nonlinear models subject to noise—to the data-driven setting. We achieve robustness by requiring that the computed set be invariant for all system dynamics compatible with the finite collected noisy measurements. Finally, leveraging the properties of DC functions, we present a tractable convex optimization algorithm that connects this condition to compute the largest possible $\lambda$-contractive set within the state constraints.
@@
Notation. The set of real numbers is denoted by $\mathbb{R}$, and the $n$-dimensional Euclidean space is $\mathbb{R}^{n}$. The state vector at time step $k$ is denoted by $x_{k}\in\mathbb{R}^{n}$, with its $j$-th component denoted by $x_{j,k}$. The operator $\nabla_{x}$ denotes the gradient with respect to the vector $x$. The infinity norm of a vector is denoted by $\|\cdot\|_{\infty}$. For every $n\in\mathbb{N}$, define $\mathbb{N}_{n}=\{i\in\mathbb{N}:1\leq i\leq n\}$. The vector of all ones with the proper dimension is denoted by $\mathbf{1}$. The interior of a set is denoted by $\text{int}(\cdot)$. The notation $\lambda\Omega$ represents the scaling of the set $\Omega$ by the constant scalar $\lambda$. Consider the matrix $H\in\mathbb{R}^{n\times n}$, the notation $H_{i}$ represents the $i$-th row of this matrix. Throughout this paper, the notation $\theta=\text{vec}(\Theta^{T})$ defines the vector $\theta$ as the vectorization of the transpose of the matrix $\Theta$. Given a set $P$, co($P$) denotes its convex hull.
II Preliminary Results
@@
In this paper, we consider the following discrete-time nonlinear system:
$x_{k+1}=f(x_{k},u_{k})+w_{k},$ (1)
@@
The sets $X$, $U$, and $W$ are assumed to be compact, convex, and to contain the origin in their interiors.
II-A Difference of Convex Functions
A key advantage of the DC framework is its generality, as a wide class of nonlinear functions can be represented as the Difference of Convex (DC) functions. This structure is particularly useful because it allows for the application of powerful tools from convex analysis to otherwise non-convex problems. Therefore, we assume the following for the system (1):
@@ II-A Difference of Convex Functions
The system dynamics vector, represented by $f(x,u):X\times U\to\mathbb{R}^{n}$, is assumed to be a vector of DC functions, differentiable at $(0,0)$. This implies there exist two functions $g_{j},h_{j}:X\times U\to\mathbb{R}$ for all $j\in\mathbb{N}_{n}$ of vector function $f(x,u)$, which are convex $\forall(x_{k},u_{k})\in X\times U$, that satisfy the following conditions:
@@
The class of DC functions is quite large: Every function with continuous second partial derivatives is a DC function on a convex compact set of its domain [9]. Consequently, the requirement in Assumption 2 that the system dynamics need to be a DC function is not a significant restriction, ensuring that the methods proposed in this paper are applicable to a broad class of nonlinear systems.
With this functional framework established, we now introduce the key definitions of $\lambda$-contractiveness and sufficient condition for a set to be contractive for nonlinear systems that are fundamental to our proposed method.
II-B $\lambda$-Contractiveness for Nonlinear Systems
Before presenting our main results, it is necessary to review the key concepts from set-theoretic control upon which our work is built. Therefore, we start with the following definitions:
@@ II-B $\lambda$-Contractiveness for Nonlinear Systems
A set $\Omega\subseteq X$ is a robust control invariant set for the system $x_{k+1}=f(x_{k},u_{k})+w_{k}$ subject to constraints $x_{k}\in X$ and $u_{k}\in U$ if, for every $x_{k}\in\Omega$, there exists a control input $u_{k}\in U$ such that $f(x_{k},u_{k})+w_{k}\in\Omega$ for all $w_{k}\in W$.
Definition 2.
[2]: Given the contraction factor $\lambda\in[0,1]$, a convex compact set $\Omega$ with $0\in\text{int}(\Omega)$ is (robust) $\lambda$-contractive for system $x_{k+1}=f(x_{k},u_{k})+w_{k}$ if for any $x_{k}\in\Omega\subseteq X$, there exists an input $u_{k}\in U$ such that $f(x_{k},u_{k})+w_{k}\in\lambda\Omega$, $\forall w_{k}\in W$.
@@
In the specific case where the set $\Omega$ is a polytope described by a system of linear inequalities, i.e., $\Omega=\{x\in\mathbb{R}^{n}:Hx\leq\textbf{1}\}$, the condition for $\lambda$-contractiveness can be stated as follows: For every $x_{k}\in\Omega\subseteq X$, there exists a control input $u_{k}\in U$ such that $\forall w_{k}\in W$, the resulting $f(x_{k},u_{k})+w_{k}$ satisfies the polytope’s defining inequalities scaled by $\lambda$: $H(f(x_{k},u_{k})+w_{k})\leq\lambda\mathbf{1}$.
Building on the definition of a robust control invariant set, we now review a key result from [8]. The main contribution of that work was to utilize the properties of DC functions to establish a sufficient condition for a set to be $\lambda$-contractive for nonlinear systems.
@@
[4]: Given a set $\Gamma\subseteq\mathbb{R}^{n}$, the support function of $\Gamma$ evaluated at $c\in\mathbb{R}^{n}$ is defined as: $\phi_{\Gamma}(c)=\sup_{x\in\Gamma}c^{T}x$.
@@
[8]: Let Assumptions 1 and 2 hold. Given the vector of DC function $f(\cdot,\cdot):X\times U\to\mathbb{R}^{n}$, whose components are decomposed as $f_{j}(x,u)=g_{j}(x,u)-h_{j}(x,u),\forall j\in\mathbb{N}_{n}$ with convex functions $g_{j}(x,u)$ and $h_{j}(x,u)$, and $c\in\mathbb{R}^{n}$, define $F(x,u,c):X\times U\times\mathbb{R}^{n}\to\mathbb{R}$ as the function
$F(x,u,c)=$ $\sum_{j\in b_{+}}c_{j}(g_{j}(x,u)-h_{j}^{L}(x,u))$ (2)
@@
The function $F(x,u,c)$ exhibits the following important property:
@@
[8]: Let Assumptions 1 and 2 hold. Given the DC function $f(\cdot,\cdot):X\times U\to\mathbb{R}^{n}$ as in (1), then for every $c\in\mathbb{R}^{n}$, the function $F(\cdot,\cdot,c)$ defined in (2) is convex in $(x,u)\in X\times U$. Moreover, it can be shown that for any $c\in\mathbb{R}^{n}$, the inequality $c^{T}f(x,u)\leq F(x,u,c)$ holds $\forall(x,u)\in X\times U$.
@@
[8]: Let Assumptions 1 and 2 hold. Consider a polytope $\Omega=\{x\in\mathbb{R}^{n}:Hx\leq\mathbf{1}\}\subseteq X$, and DC system (1) which is subject to process noise. If there exist control actions defined at the vertices $u^{p}=u_{v}^{p}\in U$, $\forall p\in\mathbb{N}_{n_{v}}$, such that
$F(v^{p},u_{v}^{p},H_{i}^{T})\leq\lambda_{\omega}-\phi_{W}(H_{i}^{T}),\quad\forall p\in\mathbb{N}_{n_{v}},\forall i\in\mathbb{N}_{n_{h}}.$ (3)
where $\lambda_{\omega}\in[0,1]$ represents the contraction factor for noisy nonlinear system, then $\Omega$ is a $\lambda_{\omega}$-contractive set for system (1) and constraints $x\in X,u\in U$.
III Problem Statement
This section formulates the data-driven synthesis of robust $\lambda$-contractive sets for the unknown nonlinear system (1). We introduce the following standing assumptions to ensure the problem is well-posed and computationally tractable:
@@ III Problem Statement
The nonlinear function $f(x_{k},u_{k})$ is assumed to lie in the span of a known dictionary of $d_{f}$ basis functions, $\phi(x,u):X\times U\to\mathbb{R}^{d_{f}}$. Consequently, there exists an unknown parameter matrix $\Theta\in\mathbb{R}^{n\times d_{f}}$ such that the system dynamics can be expressed as $f(x_{k},u_{k})=\Theta\phi(x_{k},u_{k})$.
Under these assumptions, the problem of computing the maximal robust invariant set is defined as:
@@
Under Assumptions (13), the objective is to compute the largest polytopic robust $\lambda$-contractive set, $\Omega\subseteq X$, using the finite dataset $\mathcal{D}=\left\{x_{k},u_{k},x_{k+1}\}_{k=0}^{T-1}\right$. that is generated from the noisy nonlinear system (1). The computed set must be robust $\lambda$-contractive for all possible system dynamics that are consistent with the data dictionary, $\mathcal{D}$, and the known noise upper bound $\epsilon$.
IV Data-Driven Computation of Robust Contractive Set
In this section, we address Problem 1 by developing a constructive framework for robust set synthesis. Our approach relies on a set of feasible system parameters, referred to as the consistency set, defined as follows: